# A.I.R.: Adaptive, Iterative, and Reasoning-based Frame Selection For Video Question Answering

### v1

@@ Abstract
Effectively applying Vision-Language Models (VLMs) to Video Question Answering (VideoQA) hinges on selecting a concise yet comprehensive set of frames, as processing entire videos is computationally infeasible. However, current frame selection methods face a critical trade-off: approaches relying on lightweight similarity models, such as CLIP, often fail to capture the nuances of complex queries, resulting in inaccurate similarity scores that cannot reflect the authentic query-frame relevance, which further undermines frame selection. Meanwhile, methods that leverage a VLM for deeper analysis achieve higher accuracy but incur prohibitive computational costs. To address these limitations, we propose A.I.R., a training-free approach for Adaptive, Iterative, and Reasoning-based frame selection. We leverage a powerful VLM to perform deep, semantic analysis on complex queries, and this analysis is deployed within a cost-effective iterative loop that processes only a small batch of the most high-potential frames at a time. Extensive experiments on various VideoQA benchmarks demonstrate that our approach outperforms existing frame selection methods, significantly boosts the performance of the foundation VLM, and achieves substantial gains in computational efficiency over other VLM-based techniques.
Project Page: https://ucf-air.github.io/
Codebase: https://github.com/UCF-AIR/A.I.R.
@@ 1 Introduction
Recent advancements in large Vision-Language Models (VLMs) have revolutionized tasks at the intersection of vision and text (Li et al., 2023; Hurst et al., 2024; Alayrac et al., 2022). Building on their success with static images, the focus has shifted to the more challenging domain of video understanding, which demands a model’s ability to reason over complex temporal dynamics and visual narratives. Powerful foundation VLMs are now being adapted to tackle video understanding tasks like Video Question Answering (VideoQA) (Bai et al., 2025; Zhu et al., 2025; Li et al., 2024).
@@
To alleviate the issue above, a prominent line of research focuses on query-related frame selection. Many of these approaches (Zhang et al., 2025; Sun et al., 2025; Liu et al., 2025; Tang et al., 2025) leverage lightweight multi-modal models like CLIP (Radford et al., 2021) to compute a query-frame similarity, which then guides their proposed algorithm for effective selection. However, to answer a complex query like ‘After introducing Tofu making, what kind of traditional technique or scenic spot did the youtuber introduce according to what is shown in the video?’ (Fig. 1 (a)), which requires temporal reasoning (e.g., ‘After’) and holistic semantic understanding, a lightweight model lacks these capabilities and instead treats the query as a bag of keywords. Consequently, as shown in Fig. 1 (b), it incorrectly assigns high similarity scores (score: 0.7) to frame related to the query’s context ‘Tofu’ or a visually similar frame with ‘Market’ (score: 0.8). Crucially, the frame containing the correct option ‘A Buddhist Temple’, is ranked lower (score: 0.4) because it has a weaker surface-level association with the query’s keywords. Therefore, the lightweight model produces an inaccurate similarity score that fails to reflect the true query-frame relevance, leading to the failure of the selection algorithm.
@@ 2 Related Works
Vision Language Models for Video Question Answering. The paradigm for Video Question Answering (VideoQA) has shifted from early CNN-RNN architectures (Donahue et al., 2015; Venugopalan et al., 2015) to modern Vision-Language Models (VLMs) (Alayrac et al., 2022; Li et al., 2021; Liu et al., 2023b; a). Leveraging the extensive knowledge and reasoning skills acquired during pre-training, these models have spurred two main research directions. The first focuses on models specifically adapted for the video modality, such as Video-LLaVA (Lin et al., 2023) and InternVideo (Wang et al., 2022). A second, more recent trend emphasizes general-purpose VLMs that can process both images and videos, demonstrating strong zero-shot and adaptive capabilities across various tasks. Aligning with this latter direction, our work utilizes a single, powerful VLM—selected from VILA (Lin et al., 2024), QwenVL (Bai et al., 2025; Wang et al., 2024a), LLaVA-OneVision (Li et al., 2024), and InternVL-3 (Zhu et al., 2025)—to perform both fine-grained analysis of individual frames and high-level video question answering.
@@ 3.2 Adaptive Initial Sampling
As shown in Fig. 3 (a), with the initial $n$ frames and the query, unlike other frame selection methods (Liu et al., 2025; Sun et al., 2025; Tang et al., 2025) that directly work on uniformly sampled frames with their proposed approaches, we perform an Adaptive Initial Sampling. This process selects $K$ query-related frames in advance $(K<n)$, which not only provides prior guidance, but also reduces the computation cost for our subsequent Iterative Frame Selection in Sec. 3.3. To achieve this goal, we must first adaptively separate high-relevance frames from low-relevance ones. Hence, we propose an adaptive threshold $T$, inspired by Gaussian Mixture Models (GMMs) (Huang & Chau, 2008; Zhao et al., 2019). Specifically, we hypothesize that the similarity scores $S$ are drawn from a mixture of two underlying distributions: a high-relevance cluster and a low-relevance one. We fit a GMM with two components to model these two clusters. Let the means and standard deviations of them be $(\mu_{1},\sigma_{1})$ and $(\mu_{2},\sigma_{2})$. The threshold $T$ is calculated as:
@@ 3.3 Iterative Frame Selection
For each interval, we calculate its potential—a score indicating its importance to the query—based on its corresponding slice of the pre-computed similarity signal, $S_{f_{i}:f_{i+1}}$. Inspired by signal processing principles (Boreczky & Rowe, 1996; Wolf, 1996; Liu et al., 2003), this potential is computed as a product of three factors: Relevance, Complexity, and Length (see A.2.4 for details). For a discrete similarity signal $S$, the potential of the interval $\mathrm{I}_{i}$ is calculated as:
@@
Model
LLM
Size
#Frames Video-MME MLVUdev LVBval
@@
3.4 Efficiency Analysis
$w_{\mathrm{best}}\leq n_{\mathrm{A.I.R.}}\leq w_{\mathrm{worst}},$ (7)
4 Experiments
4.1 Implementations
We use 4 widely-used foundation VLMs as our backbones: VILA-1.5-8B (Lin et al., 2024), QwenVL-2.5-7B (Bai et al., 2025), InternVL-3-8B (Zhu et al., 2025), and LLaVA-OneVision-7B (Li et al., 2024). We use the same VLM for analysis and answering. Additionally, we employ EVA-CLIP-L (Sun et al., 2023) to compute the query-frame similarity score $S$. We evaluate A.I.R. on various long video benchmarks, i.e. Video-MME (Fu et al., 2025), MLVU (Zhou et al., 2025), and LongVideoBench (LVB, Wu et al. (2024)). Besides, we also employ short video benchmarks EgoSchema (Mangalam et al., 2023) and NextQA (Xiao et al., 2021). More details are in A.3. To ensure a fair comparison with other competing frame selection methods, in Tab. 1 and 2, we re-evaluated methods with available code ($\dagger$}) in our controlled environment (lmms-eval, Zhang et al. (2024)), while for others, we compare against their reported metrics (*) in similar settings.
@@ 4.2 Comparison with the State-of-the-Art
Model
VLM
Size
#Frames Egoschema NextQA
@@ 4.3 Ablation Study
Influence of Different VLM Scales. As shown in Tab. 3, the results of various scales of VLMs reveal two key findings: (1) A.I.R. provides a consistent performance uplift across all tested scales of the QwenVL-2.5 models (7B, 32B, and 72B), confirming the robustness of our method; and (2) We observe that the accuracy gain is most pronounced on the smaller 7B model (+4.2%), while the margin narrows as the base VLM becomes more powerful. This suggests that our frame selection is most critical for smaller models with a weaker intrinsic ability to discern relevance.
@@
Experimental Efficiency Analysis (More Analysis in A.4.5). The results in Tab. 6 highlight A.I.R.’s superior efficiency-performance trade-off against competing methods on NeXTQA. When we set the worst workload $w_{\mathrm{worst}}$ to 72, which controls the max frames analyzed by VLM, our method not only achieves a SoTA accuracy of 82.6% but does so by adaptively analyzing only 32.2 frames on average. Even with a low worst workload ($w_{\mathrm{worst}}$=16), A.I.R. still attains an impressive 81.7% accuracy with just 12.4 frames, again outperforming other methods. Tab. 7 illustrates the controlled time comparison of A.I.R. against our own baseline, Direct VLM Analysis, which mimics a conventional fixed-budget manner that other methods apply using the same VLM (InternVL3-8B). For the intermediate frame analysis (VLM Analysis Time), the baseline requires 162.03s to process 128 frames, while A.I.R. takes just 42.31s by adaptively processing only 36.5 frames via Iterative Frame Selection. Beyond this VLM Analysis Time comparison, the computational overhead of our frame Adaptive Initial Sampling and Iterative Frame Selection stage is minimal, adding up to 0.21s compared to the uniform sampling baseline.
@@ 6 Reproducibility Statement
We provide the full details for reproducing our method in the main paper. All datasets used for experiments are publicly available. The partial code of our method, including the code for two core stages (i.e., Adaptive Initial Sampling and Iterative Frame Selection), will be provided in the supplementary materials. We will release the full code and model via GitHub once the paper gets accepted.
@@ A.2.4 Interval Potential Ranking
Inspired by the research in the Signal Processing domain, we define the three factors to rank an interval as: (1) Relevance: The average signal magnitude, analogous to the DC component (Oppenheim, 1999). Higher values suggest greater pertinence; (2) Complexity: The signal’s Total Variation, which quantifies volatility and often highlights significant events like scene changes (Rudin et al., 1992); (3) Length: The logarithmic duration of the interval, which prioritizes larger unexplored segments while modeling diminishing returns (Pirolli & Card, 1999). Thus, the potential of interval $\mathrm{I}_{i}$ is calculated as:
@@ A.3.2 Experimental Environment
All of our experiments are conducted using NVIDIA GH200 GPUs, with Arch64 as the CPU architecture. Our inference is using Pytorch 2.5 with CUDA-11.6 and GCC-11.4.0. One-time experiment of our method is run using 4 aforementioned machines via lmms-eval (Zhang et al., 2024) inference framework. When using InterVL3 as our foundation VLM, the inference time (Zhu et al., 2025) is around 10 hours. For Video-MME w/o subtitle (Fu et al., 2025). The cost time ranges from 9 hours to 17 hours per different VLMs. Due to the limit of lmms-eval framework, all of our inference is under 1 batch size, while we set the sub-batch size to 12 (e.g., for VLM analysis on frames). For CLIP cache, because each benchmark has queries using the same video repeatedly, we store the image features for quick cosine-similarity computation. All of our videos are loaded from high-res files (all data are provided by lmms-eval, https://github.com/EvolvingLMMs-Lab/lmms-eval), so preprocessing operations like Resizing may increase the time for computation of CLIP features.
@@ A.4.1 Additional Comparison with SoTAs
Besides the main comparison with SoTAs using QwenVL-2.5, we conducted an additional set of experiments with Qwen2-VL-7B (Wang et al., 2024a) (Tab. 8) to further demonstrate the generalizability and robustness of our A.I.R. framework. We compare our performance against a 32-frame uniform sampling baseline as well as several other state-of-the-art (SoTA) frame selection methods. The results are presented in Tab. 8. The data shows that A.I.R. consistently outperforms both the baseline and competing methods on this new VLM backbone. Notably, on the Video-MME (w/o sub) benchmark, our approach achieves the highest accuracy of 60.0%, surpassing strong methods like AKS (59.9%) and Q-Frame (58.3%), and providing a significant +2.4% uplift over the baseline. Our method also achieves the top performance on Egoschema and NextQA. While AKS shows a strong result on LVB, our method remains highly competitive with a score of 58.9%, a substantial improvement over the baseline’s 55.5%. This experiment validates that the benefits of our adaptive, iterative, and reasoning-based approach are not tied to a specific VLM architecture and can be effectively generalized.

### v2

@@ Abstract
Effectively applying Vision-Language Models (VLMs) to Video Question Answering (VideoQA) hinges on selecting a concise yet comprehensive set of frames, as processing entire videos is computationally infeasible. However, current frame selection methods face a critical trade-off: approaches relying on lightweight similarity models, such as CLIP, often fail to capture the nuances of complex queries, resulting in inaccurate similarity scores that cannot reflect the authentic query-frame relevance, which further undermines frame selection. Meanwhile, methods that leverage a VLM for deeper analysis achieve higher accuracy but incur prohibitive computational costs. To address these limitations, we propose A.I.R., a training-free approach for Adaptive, Iterative, and Reasoning-based frame selection. We leverage a powerful VLM to perform deep, semantic analysis on complex queries, and this analysis is deployed within a cost-effective iterative loop that processes only a small batch of the most high-potential frames at a time. Extensive experiments on various VideoQA benchmarks demonstrate that our approach outperforms existing frame selection methods, significantly boosts the performance of the foundation VLM, and achieves substantial gains in computational efficiency over other VLM-based techniques. Project Page: https://ucf-air.github.io/
@@ 1 Introduction
Recent advancements in large Vision-Language Models (VLMs) have revolutionized tasks at the intersection of vision and text (Li et al., 2023; Hurst et al., 2024; Alayrac et al., 2022). Building on their success with static images, the focus has shifted to the more challenging domain of video understanding, which demands a model’s ability to reason over complex temporal dynamics and visual narratives. Powerful foundation VLMs are now being adapted to tackle video understanding tasks like Video Question Answering (VideoQA) (Bai et al., 2025; Zhu et al., 2025; Li et al., 2024a).
@@
To alleviate the issue above, a prominent line of research focuses on query-related frame selection. Many of these approaches (Zhang et al., 2025; Sun et al., 2025; Liu et al., 2025; Tang et al., 2025) leverage lightweight multi-modal models like CLIP (Radford et al., 2021) to compute a query-frame similarity, which then guides their proposed algorithm for effective selection. However, to answer a complex query like ‘After introducing Tofu making, what kind of traditional technique or scenic spot did the youtuber introduce according to what is shown in the video?’ (Fig. 1 (a)), which requires temporal reasoning (e.g., ‘After’) and holistic semantic understanding, a lightweight model lacks these capabilities and instead treats the query as a bag of keywords Yuksekgonul et al. (2022); Xie et al. (2025). Consequently, as shown in Fig. 1 (b), it incorrectly assigns high similarity scores (score: 0.7) to frame related to the query’s context ‘Tofu’ or a visually similar frame with ‘Market’ (score: 0.8). Crucially, the frame containing the correct option ‘A Buddhist Temple’, is ranked lower (score: 0.4) because it has a weaker surface-level association with the query’s keywords. Therefore, the lightweight model produces an inaccurate similarity score that fails to reflect the true query-frame relevance, leading to the failure of the selection algorithm.
@@ 2 Related Works
Vision Language Models for Video Question Answering. The paradigm for Video Question Answering (VideoQA) has shifted from early CNN-RNN architectures (Donahue et al., 2015; Venugopalan et al., 2015) to modern Vision-Language Models (VLMs) (Alayrac et al., 2022; Li et al., 2021; Liu et al., 2023b; a). Leveraging the extensive knowledge and reasoning skills acquired during pre-training, these models have spurred two main research directions. The first focuses on models specifically adapted for the video modality, such as Video-LLaVA (Lin et al., 2023) and InternVideo (Wang et al., 2022). A second, more recent trend emphasizes general-purpose VLMs that can process both images and videos, demonstrating strong zero-shot and adaptive capabilities across various tasks. Aligning with this latter direction, our work utilizes a single, powerful VLM—selected from VILA (Lin et al., 2024), QwenVL (Bai et al., 2025; Wang et al., 2024a), LLaVA-OneVision (Li et al., 2024a), and InternVL-3 (Zhu et al., 2025)—to perform both fine-grained analysis of individual frames and high-level video question answering.
@@ 3.2 Adaptive Initial Sampling
As shown in Fig. 3 (a), with the initial $n$ frames and the query, unlike other frame selection methods (Liu et al., 2025; Sun et al., 2025; Tang et al., 2025) that directly work on uniformly sampled frames with their proposed approaches, we perform an Adaptive Initial Sampling. This process selects $K$ query-related frames in advance $(K<n)$, which not only provides prior guidance, but also reduces the computation cost for our subsequent Iterative Frame Selection in Sec. 3.3. To achieve this goal, we must first adaptively separate high-relevance frames from low-relevance ones. Hence, we propose an adaptive threshold $T$, inspired by Gaussian Mixture Models (GMMs) (Huang and Chau, 2008; Zhao et al., 2019). Specifically, we hypothesize that the similarity scores $S$ are drawn from a mixture of two underlying distributions: a high-relevance cluster and a low-relevance one. We fit a GMM with two components to model these two clusters. Let the means and standard deviations of them be $(\mu_{1},\sigma_{1})$ and $(\mu_{2},\sigma_{2})$. The threshold $T$ is calculated as:
@@ 3.3 Iterative Frame Selection
For each interval, we calculate its potential—a score indicating its importance to the query—based on its corresponding slice of the pre-computed similarity signal, $S_{f_{i}:f_{i+1}}$. Inspired by signal processing principles (Boreczky and Rowe, 1996; Wolf, 1996; Liu et al., 2003), this potential is computed as a product of three factors: Relevance, Complexity, and Length (see A.2.4 for details). For a discrete similarity signal $S$, the potential of the interval $\mathrm{I}_{i}$ is calculated as:
@@
3.4 Efficiency Analysis
$w_{\mathrm{best}}\leq n_{\mathrm{A.I.R.}}\leq w_{\mathrm{worst}},$ (7)
4 Experiments
4.1 Implementations
We use 4 widely-used foundation VLMs as our backbones: VILA-1.5-8B (Lin et al., 2024), QwenVL-2.5-7B (Bai et al., 2025), InternVL-3-8B (Zhu et al., 2025), and LLaVA-OneVision-7B (Li et al., 2024a). We use the same VLM for analysis and answering. Additionally, we employ EVA-CLIP-L (Sun et al., 2023) to compute the query-frame similarity score $S$. We evaluate A.I.R. on various long video benchmarks, i.e. Video-MME (Fu et al., 2025), MLVU (Zhou et al., 2025), and LongVideoBench (LVB, Wu et al. (2024)). Besides, we also employ short video benchmarks EgoSchema (Mangalam et al., 2023) and NextQA (Xiao et al., 2021). More details are in A.3. To ensure a fair comparison with other competing frame selection methods, in Tab. 1 and 2, we re-evaluated methods with available code ($\dagger$}) in our controlled environment (lmms-eval, Zhang et al. (2024a)), while for others, we compare against their reported metrics (*) in similar settings.
@@ 4.1 Implementations
Model LLM Size #Frames Video-MME MLVUdev LVBval
@@ 4.2 Comparison with the State-of-the-Art
Model VLM Size #Frames Egoschema NextQA
@@ 4.3 Ablation Study
Influence of Different VLM Scales. As shown in Tab. 3, the results of various scales of VLMs reveal two key findings: (1) A.I.R. provides a consistent performance uplift across all tested scales of the QwenVL-2.5 models (7B, 32B, and 72B), confirming the robustness of our method; and (2) We observe that the accuracy gain is most pronounced on the smaller 7B model (+4.2%), while the margin narrows as the base VLM becomes more powerful. This suggests that our frame selection is most critical for smaller models with a weaker intrinsic ability to discern relevance. We further validate the scalability of our plug-and-play framework on state-of-the-art models like GPT-4o, demonstrating consistent performance gains even when scaling to extremely large VLMs.
@@
Experimental Efficiency Analysis (More Analysis in A.4.5).
The results in Tab. 6 highlight A.I.R.’s superior efficiency-performance trade-off against competing methods on NextQA. When we set the worst workload $w_{\mathrm{worst}}$ to 72, which controls the max frames analyzed by VLM, our method not only achieves a SoTA accuracy of 82.6% but does so by adaptively analyzing only 32.2 frames on average. Even with a low worst workload ($w_{\mathrm{worst}}$=16), A.I.R. still attains an impressive 81.7% accuracy with just 12.4 frames, again outperforming other methods. Tab. 7 further demonstrates A.I.R.’s computational efficiency on Video-MME. Tab. 7 illustrates the controlled time comparison of A.I.R. against our own baseline, Direct VLM Analysis, which mimics a conventional fixed-budget manner that other methods apply using the same VLM (InternVL3-8B). For the intermediate frame analysis (VLM Analysis Time), the baseline requires 162.03s to process 128 frames, while A.I.R. takes just 42.31s by adaptively processing only 36.5 frames via Iterative Frame Selection. Beyond this VLM Analysis Time comparison, the computational overhead of our frame Adaptive Initial Sampling and Iterative Frame Selection stage is minimal, adding up to 0.21s compared to the uniform sampling baseline.
Table 8: Generalization results on temporal grounding benchmark Charades-STA (Gao et al., 2017).
Model R1@0.3 R1@0.5 R1@0.7 mIoU
Trained Temporal Grounding VideoLLMs
VTimeLLM (Huang et al., 2024) 51.0 27.5 11.4 31.2
HawkEye (Wang et al., 2024c) 50.6 31.4 14.5 33.7
TimeChat (Ren et al., 2024) - 32.2 13.4 30.6
TimeSuite (Zeng et al., 2024) 69.9 48.7 24.0 -
TRACE (Guo et al., 2024) - 40.3 19.4 -
General VideoLLMs
GPT-4o (Hurst et al., 2024) 55.0 32.0 11.5 35.4
Qwen2.5-VL-7B (Bai et al., 2025) 44.5 30.3 15.2 30.1
LongVA-7B-DPO (Zhang et al., 2024b) 22.6 10.1 2.2 14.6
Aria (Li et al., 2024b) 39.0 18.6 6.6 26.7
Trained Frame Selection Methods
GenS (Yao et al., 2025) 62.9 38.7 15.2 38.0
Training-Free Frame Selection Methods
[…]
Generalization to Grounding Task Analysis. To demonstrate that A.I.R.’s adaptive frame selection generalizes beyond VideoQA, we evaluate its performance on Charades-STA (Gao et al., 2017), a challenging temporal grounding benchmark where models must localize specific moments in videos based on natural language queries. As shown in Tab. 8, our training-free method achieves strong performance across all metrics: 59.5% R1@0.3, 39.5% R1@0.5, 18.0% R1@0.7, and 38.8% mIoU. Notably, A.I.R. substantially outperforms general-purpose VideoLLMs like Qwen2.5-VL-7B (44.5% $\rightarrow$ 59.5% R1@0.3) and even surpasses GPT-4o (32.0% vs. 39.5% R1@0.5). More impressively, our method achieves comparable or superior results to GenS (Yao et al., 2025), a trained frame selection method specifically designed for grounding tasks, while remaining completely training-free. While specialized temporal grounding models like TimeSuite (Zeng et al., 2024) achieve higher performance through task-specific training, our results demonstrate that A.I.R.’s adaptive sampling and iterative refinement effectively identify temporally relevant segments without requiring any grounding-specific supervision. This validates the broad applicability of our approach across diverse video understanding tasks.
@@ 6 Reproducibility Statement
We provide the full details for reproducing our method in the main paper. All datasets used for experiments are publicly available. The partial code of our method, including the code for two core stages (i.e., Adaptive Initial Sampling and Iterative Frame Selection), will be provided in the supplementary materials.
Acknowledgments
This work used Delta at UIUC NCSA through allocation CIS250367 and 250473 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S. NSF grants 2138259, 2138286, 2138307, 2137603, and 2138296.
@@ A.2.4 Interval Potential Ranking
Inspired by the research in the Signal Processing domain, we define the three factors to rank an interval as: (1) Relevance: The average signal magnitude, analogous to the DC component (Oppenheim, 1999). Higher values suggest greater pertinence; (2) Complexity: The signal’s Total Variation, which quantifies volatility and often highlights significant events like scene changes (Rudin et al., 1992); (3) Length: The logarithmic duration of the interval, which prioritizes larger unexplored segments while modeling diminishing returns (Pirolli and Card, 1999). Thus, the potential of interval $\mathrm{I}_{i}$ is calculated as:
@@ A.3.2 Experimental Environment
All of our experiments are conducted using NVIDIA GH200 GPUs, with Arch64 as the CPU architecture. Our inference is using Pytorch 2.5 with CUDA-11.6 and GCC-11.4.0. One-time experiment of our method is run using 4 aforementioned machines via lmms-eval (Zhang et al., 2024a) inference framework. When using InterVL3 as our foundation VLM, the inference time (Zhu et al., 2025) is around 10 hours. For Video-MME w/o subtitle (Fu et al., 2025). The cost time ranges from 9 hours to 17 hours per different VLMs. Due to the limit of lmms-eval framework, all of our inference is under 1 batch size, while we set the sub-batch size to 12 (e.g., for VLM analysis on frames). For CLIP cache, because each benchmark has queries using the same video repeatedly, we store the image features for quick cosine-similarity computation. All of our videos are loaded from high-res files (all data are provided by lmms-eval, https://github.com/EvolvingLMMs-Lab/lmms-eval), so preprocessing operations like Resizing may increase the time for computation of CLIP features.
@@ A.4.1 Additional Comparison with SoTAs
Comparison with other SOTA Methods on Qwen2-VL-7B. Besides the main comparison with SoTAs using QwenVL-2.5, we conducted an additional set of experiments with Qwen2-VL-7B (Wang et al., 2024a) (Tab. 9) to further demonstrate the generalizability and robustness of our A.I.R. framework. We compare our performance against a 32-frame uniform sampling baseline as well as several other state-of-the-art (SoTA) frame selection methods. The results are presented in Tab. 9. The data shows that A.I.R. consistently outperforms both the baseline and competing methods on this new VLM backbone. Notably, on the Video-MME (w/o sub) benchmark, our approach achieves the highest accuracy of 60.0%, surpassing strong methods like AKS (59.9%) and Q-Frame (58.3%), and providing a significant +2.4% uplift over the baseline. Our method also achieves the top performance on Egoschema and NextQA. While AKS shows a strong result on LVB, our method remains highly competitive with a score of 58.9%, a substantial improvement over the baseline’s 55.5%. This experiment validates that the benefits of our adaptive, iterative, and reasoning-based approach are not tied to a specific VLM architecture and can be effectively generalized.
[…]
Table 10: Comparison of Agent-based VLM Frame Selection Methods on Video Question Answering Benchmarks.
Method Venue Base Model #Frames Video-MME MLVU LVB EgoSchema NextQA
w/o sub w/ sub Full Subset
Agent-based Frame Selection Methods
VideoLucy (Zuo et al., ) NeurIPS’25 DeepSeek-R1 - 72.5 - 76.1 - - - -
VideoRAG (Luo et al., 2024) NeurIPS’25 LLaVA-Video-7B 64 - - 72.4 58.7 - - -
TYe et al. (2025) CVPR’25 LLaVA-OV-7B 8 - - - - - 66.6 80.4
DrVideo (Ma et al., 2025) CVPR’25 GPT-4 - - - - - 61.0 66.4 -
MemVid (Yuan et al., 2025) arXiv’25 Qwen2VL-7B 128 63.7 65.7 - - - - -
VideoAgent2 (Zhi et al., 2025) arXiv’25 GPT-4o - - - - - 75.4 - 80.5
AKEYS (Fan et al., 2025) arXiv’25 GPT-4o $\leq$32 - - - - 63.6 68.6 78.1
Our Frame Selection Method for Comparison
[…]